OpenStudy (anonymous):

What are the optimized dimensions of a cylinder that would result in a maximum volume given a surface area of 450 cm squared Please Explain =]

OpenStudy (anonymous):

volume = \[\pi r^2h\] yes

OpenStudy (amistre64):

and surface area = 2pi r^2 + 2pi rh

OpenStudy (amistre64):

use the surface area to determine the value of one variable in terms of the other; and derive the volume

OpenStudy (anonymous):

surface area is \[4\pi r^2+2\pi r h=450\]

OpenStudy (amistre64):

2(p r^2) = 2pi r^2 maybe?

OpenStudy (anonymous):

solve second one for h, plug into first one to get an equation in one varaible

OpenStudy (anonymous):

oh of course you are right. it is \[2\pi r^2+2\pi r h\]

OpenStudy (anonymous):

back to the dungeon for me

OpenStudy (amistre64):

and take the mop with you :)

OpenStudy (anonymous):

lets see if i can do next part right \[h=\frac{450-2 \pi r^2}{2 \pi r}=\frac{225}{\pi r}-r\]

OpenStudy (anonymous):

making \[V(r)=\pi r^2 (\frac{225}{\pi r}-r)\]

OpenStudy (anonymous):

\[V(r)=225r-\pi r^3\]

OpenStudy (anonymous):

take the derivative, set = 0, solve for r and be done

OpenStudy (anonymous):

i get \[V'(r)=225-3\pi r^2\]

OpenStudy (anonymous):

put \[225-3 \pi r^2=0\] \[r^2=\frac{225}{3\pi}=\frac{75}{\pi}\]

OpenStudy (anonymous):

and therefore \[r=\frac{5\sqrt{3}}{\sqrt{\pi}}\]

OpenStudy (anonymous):

is your max

OpenStudy (anonymous):

i mean it will give you the max volume

OpenStudy (anonymous):

cool thnx

OpenStudy (anonymous):

yw

OpenStudy (amistre64):

that answer seems a bit contrived dontch think ;)

OpenStudy (anonymous):

i was wondering well i hve the answer just not the wrking itss bh=9.6 cm & r= 4.8 cm

OpenStudy (amistre64):

then its good :)

OpenStudy (anonymous):

*its not bh it h=9.6 cm

OpenStudy (anonymous):

kk gracias

OpenStudy (anonymous):

thhnxx again